OpenStudy (beccab003):

Don't understand what discriminant's are. Help. Medal to best answer. Find the discriminant of the quadratic equation 3x^2 – x + 5 = 0

OpenStudy (beccab003):

my answer is: -59, no real roots possible. I just plugged this into my graphing calculator and that's where I got my answer because it doesn't cross the x axis. Am I right?

OpenStudy (beccab003):

But I still don't know what discriminant's are...@Directrix @jim_thompson5910 @DanJS

OpenStudy (dtan5457):

Using that formula, you can find out what type of roots your gonna have

OpenStudy (zale101):

Here, you have the quadratic equation 3x^2 - x + 5 = 0 written in a standard form of ax^2+bc+c=0; that means: a =3, b=-1, and c=5

OpenStudy (zale101):

plug a =3, b=-1, and c=5 into the discriminate and you're done! :)

OpenStudy (beccab003):

Okay. Thanks. So I was right?

OpenStudy (zale101):

Yes, descrimintate is \(b^2-4ac\) from the quadratic formula, and if you plug in the values in, \((-1)^2-4(3)(5)=-59\) Correct! :)

OpenStudy (zale101):

discriminant*

OpenStudy (beccab003):

Thanks!

OpenStudy (zale101):

No problem!